# Improving noisy free-energy measurements by adding more noise

### v2

@@ III.2 Information erasure
In this section we consider a nonlinear protocol in which a double-well potential is changed into a single-well one, a form of information erasure, with a consequent increase in the free energy of the system. The potential appearing in Eqs. (3) and (13) is now
@@
As before, increasing $\lambda$ also allows us to estimate $\beta\Delta F$ accurately, even for the short trajectory length $t_{\rm f}=1$. The energy barrier separating the initial double wells is $c_{1}(0)^{2}/2=12.5$, in units of $k_{\rm B}T$, and so the free-energy difference between initial and final forms of the potential is, to a very good approximation, $\beta\Delta F=\ln 2$. This result is closely related to the Landauer observation that erasing one bit of information at temperature $T$ costs at least $k_{\rm B}T\ln 2$ units of work Landauer (1961); Bennett and Landauer (1985). We show, in panels (c) and (d) of Fig. 5, that as $\lambda$ increases the free-energy estimator $J_{\lambda}$ tends to $\ln 2$ and the block-average variance $\sigma_{B}^{2}$ diminishes. The faster relaxation engineered by the added noise allows the particle to remain close to equilibrium throughout the transformation, reducing the dissipated work and the fluctuations associated with the Jarzynksi free-energy estimator.
@@ IV Conclusions
Here we have proposed the alternative and counterintuitive strategy of adding noise to the system in order to enhance the precision of free-energy estimates. By introducing additional stochastic fluctuations and rescaling the system’s potential energy, we can increase the system’s relaxation rate without changing its thermodynamics, and so reduce the magnitude of work fluctuations and suppress rare-event sampling issues. Using numerical simulations of model systems, we have demonstrated that this approach significantly improves the accuracy of free-energy estimates.
However, the practical applicability of this method is limited, because it depends on our ability to effect a uniform scaling of the potential energy landscape. This is not feasible in many settings, such as experiments involving single biomolecules. We can exert some control over a biomolecule’s potential energy landscape, such as by adding NaCl to a solution containing DNA, but not to the required degree. Systems involving rigid particles in optical traps Mart nez et al. (2012); Chupeau et al. (2018); Saha et al. (2023) and thermodynamic computers Aifer et al. (2024) do satisfy our requirements: external noise can be added to these systems, and the potential energy scaled in the required way. For instance, the method could be used with thermodynamic computers Conte et al. (2019); Aifer et al. (2024) as a way to accurately calculate the integrals in (5), with $U({\bm{x}},{\bm{c}}(0))$ corresponding to a simple reference state and ${\rm e}^{-\beta U({\bm{x}},{\bm{c}}(t_{\rm f}))}$ the desired integrand. However, for particle-in-trap systems we usually know the free energies of the relevant states, making the method redundant (in this case the method could be used to simply enact a given protocol more quickly Whitelam (2024)).
In general terms, our results indicate that noise engineering could be a valuable tool for enhancing thermodynamic measurements and computations, and suggest that the notion of optimal control could be broadened to include control of noise, which is now experimentally feasible, as well as control of the potential.
@@ V Acknowledgments
The convergence properties of the Jarzynski equality, Eq. (3), can be understood in a general sense by assuming that the work distribution $P_{\rm F}(W)$ is Gaussian with mean $\langle W\rangle$ and variance $\sigma^{2}$. In this case the expectation in (3) can be written
@@
Two things are apparent from Eq. (A). First, in order to evaluate the integral by trajectory sampling, we must collect good statistics at the atypical work value $W_{0}=\langle W\rangle-\sigma^{2}\beta$ that dominates the integral. This value occurs with probability $P_{\rm F}(W_{0})\sim{\rm e}^{-\sigma^{2}\beta^{2}/2}$, meaning that for good sampling of (3) we must generate a number of trajectories $N_{\rm traj}\sim P_{\rm F}(W_{0})^{-1}$ that increases exponentially with the variance $\sigma^{2}$. Second, by equating the right-hand side of (3) with the right-hand side of (A) (noting that the second line is unity), we see that

### v3

@@ III.2 Information erasure
In this section we consider a nonlinear protocol in which a double-well potential is changed into a single-well one, a form of information erasure, with a consequent increase in the free energy of the system. The work statistics $P(W)$ for this system is not in general Gaussian. The potential appearing in Eqs. (3) and (15) is now
@@
As before, increasing $\lambda$ also allows us to estimate $\beta\Delta F$ accurately, even for the short trajectory length $t_{\rm f}=1$. The energy barrier separating the initial double wells is $c_{1}(0)^{2}/2=12.5$, in units of $k_{\rm B}T$, and so the free-energy difference between initial and final forms of the potential is, to a very good approximation, $\beta\Delta F=\ln 2$. This result is closely related to the Landauer observation that erasing one bit of information at temperature $T$ costs at least $k_{\rm B}T\ln 2$ units of work Landauer (1961); Bennett and Landauer (1985). We show, in panels (c) and (d) of Fig. 5, that as $\lambda$ increases the free-energy estimator $J_{\lambda}$ tends to $\ln 2$ and the block-average variance $\sigma_{B}^{2}$ diminishes. The faster relaxation engineered by the added noise allows the particle to remain close to equilibrium throughout the transformation, reducing the dissipated reduced work and the fluctuations associated with the Jarzynksi free-energy estimator.
@@ IV Conclusions
Here we have proposed the alternative and counterintuitive strategy of adding noise to the system in order to enhance the precision of free-energy estimates. By introducing additional stochastic fluctuations and rescaling the system’s potential energy accordingly, we can increase the system’s relaxation rate without changing its thermodynamics, and so reduce the magnitude of work fluctuations and suppress rare-event sampling issues. Using numerical simulations of two model systems, we have demonstrated that this approach significantly improves the accuracy of free-energy estimates.
However, the practical applicability of this method is limited, because it depends on our ability to effect a uniform scaling of the potential energy landscape. This is not feasible in many settings, such as experiments involving single biomolecules. We can exert some control over a biomolecule’s potential energy landscape, such as by adding NaCl to a solution containing DNA, but not to the required degree. Systems involving rigid particles in optical traps Martinez et al. (2012); Chupeau et al. (2018); Saha et al. (2023) and thermodynamic computers Aifer et al. (2024) do satisfy our requirements: external noise can be added to these systems, and the potential energy scaled in the required way. For instance, the method could be used with thermodynamic computers Conte et al. (2019); Aifer et al. (2024) as a way to accurately calculate the integrals in (5), with $U({\bm{x}},{\bm{c}}(0))$ corresponding to a simple reference state and ${\rm e}^{-\beta U({\bm{x}},{\bm{c}}(t_{\rm f}))}$ the desired integrand. However, for particle-in-trap systems we usually know the free energies of the relevant states, making the method redundant (in this case the method could be used to simply enact a given protocol more quickly Whitelam (2024)).
The same time rescaling could in principle be done for underdamped Langevin dynamics, although the experimental requirements (beyond adding noise) are even more difficult to achieve than in the overdamped case. The underdamped Langevin equation reads
$m\ddot{x}_{i}+\gamma\dot{x}_{i}=-\frac{\partial U({\bm{x}})}{\partial x_{i}}+\sqrt{2\gamma k_{\rm B}T}\,\eta_{i}(t),$ (25)
where first term represents inertial effects ($m$ is mass), the second term is the damping force ($\gamma=1/\mu$ is the damping coefficient), and the noise term models thermal fluctuations. Rescaling time by a factor $\lambda\geq 1$ in (25) results in the equation
$m\ddot{x}_{i}+\lambda\gamma\dot{x}_{i}=-\lambda^{2}\frac{\partial U({\bm{x}})}{\partial x_{i}}+\sqrt{2\lambda^{3}\gamma k_{\rm B}T}\,\eta_{i}(t).$ (26)
We can produce (26) from (25) – and so effectively increase the time constant of the system – by adding to (25) a Gaussian white noise with zero mean and variance $2(\lambda^{3}-1)\gamma k_{\rm B}T$, and rescaling the potential $V$ by a factor of $\lambda^{2}$. However, we also need to rescale the damping term by a factor of $\lambda$, which in an optical trap setup would require imposing additional frictional forces.
In general terms, our results indicate that noise engineering could be a useful tool for enhancing thermodynamic measurements and computations, and suggest that the notion of optimal control could be broadened to include control of noise, which is now experimentally feasible, as well as control of the potential.
@@ V Acknowledgments
The convergence properties of the Jarzynski equality, Eq. (3), can be illustrated by assuming that the work distribution $P_{\rm F}(W)$ is Gaussian with mean $\langle W\rangle$ and variance $\sigma^{2}$. In this case the expectation in (3) can be written
@@
Two things are apparent from Eq. (A). First, in order to evaluate the integral by trajectory sampling, we must collect good statistics at the atypical work value $W_{0}=\langle W\rangle-\sigma^{2}\beta$ that dominates the integral. This value occurs with probability $P_{\rm F}(W_{0})\sim{\rm e}^{-\sigma^{2}\beta^{2}/2}$, meaning that for good sampling of (3) we must generate a number of trajectories $N_{\rm traj}\sim P_{\rm F}(W_{0})^{-1}$ that increases exponentially with the variance $\sigma^{2}$ of work fluctuations. Second, by equating the right-hand side of (3) with the right-hand side of (A) (noting that the second line is unity), we see that